# ioPUF+: A PUF Based on I/O Pull-Up/Down Resistors for Secret Key Generation in IoT Nodes

### v1

@@ Abstract
In this work, we present ioPUF+, which incorporates a novel Physical Unclonable Function (PUF) that generates unique fingerprints for Integrated Circuits (ICs) and the IoT nodes encompassing them. The proposed PUF generates device-specific responses by measuring the pull-up and pull-down resistor values on the I/O pins of the ICs, which naturally vary across chips due to manufacturing-induced process variations. Since these resistors are already integrated into the I/O structures of most ICs, ioPUF+ requires no custom circuitry, and no new IC fabrication. This makes ioPUF+ suitable for cost-sensitive embedded systems built from Commercial Off-The-Shelf (COTS) components. Beyond introducing a new PUF, ioPUF+ includes a complete datapath for converting raw PUF responses into cryptographically usable secret keys using BCH error correction and SHA-256 hashing. Further ioPUF+ also demonstrate a practical use case of PUF derive secret keys in securing device-to-device communication using AES-encryption.
We implemented ioPUF+ on the Infineon PSoC-5 microcontroller and evaluated its performance across 30 devices using standard PUF metrics. The results show excellent reliability (intra-device Hamming distance of 100.00%), strong uniqueness (inter-device Hamming distance of 50.33%), near-ideal uniformity (50.54%), and negligible bit aliasing. Stability tests under temperature and supply-voltage variations show worst-case bit-error rates of only 2.63% and 2.10%, respectively. We also profiled the resource and energy usage of the complete ioPUF+ system, including the PUF primitive, BCH decoding, SHA-256 hashing, and AES encryption. The full implementation requires only 19.8 KB of Flash, exhibits a latency of 600 ms, and consumes 79 mW of power, demonstrating the suitabilitiy of ioPUF+ for resource-constrained IoT nodes.
@@ I Introduction
In many Internet of Things (IoT) applications, securing data is paramount. Consider the case of wearable sensors that continuously collect and transmit sensitive health information from a patient to a nearby hub over the air. If the transmitted data is not adequately protected, an adversary in the vicinity may eavesdrop on the wireless channel, gain access to private health information, and potentially misuse it [2].
@@
A common approach to secure data in transit or at rest is to encrypt it using a secret key, often stored in non-volatile memory of the device, following well-established standards such as AES. However, this method has a critical limitation: if an attacker gains physical access to the device, the stored key can be extracted [3]. Once compromised, the adversary can eavesdrop on future encrypted communications or decrypt any stored information.
Public-key cryptography offers a solution, in particular for protecting data during transit. Here, the recipient’s public key is provisioned to the IoT node during an enrollment phase. Later, the node encrypts its data using the public key, before transmitting over the air. Only the intended recipient, who owns the corresponding private key, can decrypt it. However, this method has two major drawbacks. First, public-key cryptography incurs significantly higher computational and energy overheads compared to symmetric-key methods like AES, making it less suitable for resource and energy constrained IoT nodes [4]. Second, if an attacker learns the data format, they may create a malicious node that encrypts forged data using the same public key. The receiver may then process these malicious data as legitimate, leading to data-integrity violations [3]. In safety-critical systems such as in industrial IoT systems, such violations may trigger catastrophic failures or prolonged downtime.
Another possibility is to protect unauthorized readouts from device’s memory storing the key. The method works by disabling the the readout circuit to the memory once the key is logged in, so that external devices e.g., programmers cannot perform any memory reads from the device. However, the method cannot prevent sophisticated invasive attacks such as the case where the chip is decapsulated and the internal circuitry is probed to infer the key.
@@
As an alternative to the aforementioned traditional approaches, Physically Unclonable Function (PUF)-based methods for securing IoT nodes have been proposed. PUF-based techniques eliminate the need to store keys in non-volatile memory [6, 7, 8], thereby reducing the risk of key extraction and the overhead of protecting stored keys. Instead, they generate secret keys during the runtime by exploiting intrinsic hardware characteristics unique to each device, making these keys extremely difficult to duplicate. The generated key resides only in volatile memory (e.g., RAM). As such, when the device powers off, the key disappears from the memory, preventing an attacker from extracting it through physical access. Furthermore, because the key exists only during the powered-on state, additional layers of protection, such as active tamper-detection sensors or probe detection circuits, which operate only when the device is powered, can be employed.
@@
At the core of any PUF-based method for generating secret keys lies a PUF primitive, a circuit and/or algorithm that produces unique, device-specific responses that serve as a fingerprint of the hardware. Several types of PUF primitives have been reported in the literature. However, many of them suffer from a significant drawback: they require custom circuitry and, consequently, a dedicated Integrated Circuit (IC) [9]. For example, the Ring Oscillator PUF (RO-PUF) [10] relies on a specially engineered ring-oscillator network. As such many of the existing PUF primitives [11, 12, 13] have limited application in IoT nodes, which are typically built using low-cost, resource-constrained Commercial Off-The-Shelf (COTS) components. Memory-based PUFs, such as SRAM PUFs [14], are compatible with COTS microcontrollers but introduce their own constraints: response generation is typically restricted to the device’s boot phase. As a result, the derived secret key must be stored in memory throughout the device’s runtime, which may raise security concerns.
@@
In response to these limitations, we introduce ioPUF+, which incorporates a novel PUF primitive that generates unique fingerprints for Integrated Circuits (ICs), specifically, microcontroller ICs commonly used in IoT nodes, by measuring the values of the pull-up and pull-down resistors connected to the IC’s input–output (IO) pins. Our approach is based on the observation that, in many IC designs, the pull-up/down resistor values exhibit substantial inter-device variation while showing much smaller intra-device variation across repeated measurements. By leveraging these resistor values, the proposed PUF primitive produces device-specific responses without requiring custom circuitry or new IC fabrication, as pull-up and pull-down resistors are already integral to the IO structures of most COTS microcontroller ICs. Moreover, unlike SRAM PUFs, which restrict response generation to the boot phase, the responses here can be generated at any point during device operation. Consequently, ioPUF+ based secret keys can be derived on need basis and can be securely erased from memory after use, improving security of the keys.
@@
The rest of the paper is organized as follows: Section II reviews related work on I/O-based PUFs and PUFs for secret key generation used in cryptography applications. Section III presents the system design of our ioPUF+, the algorithm used for generating ioPUF+ responses, and the procedure for obtaining the PUF ID and secret key from the raw ioPUF+ response. Section IV describes the experimental setup, including hardware implementation on the PSoC5 microcontroller, evaluation of standard PUF metrics, and discussion of ioPUF+ results for these metrics. In Section V, we demonstrate cryptographic key generation, provide an end-to-end secure communication application, and discuss resource utilization for all components on the PSoC5 microcontroller. Section VI concludes the paper with future directions.
@@ II Related Work
However, ioPUF [1] design had several shortcomings. First, it measured either the pull-up or the pull-down resistor of each I/O pin, limiting the PUF response length to 45 bits when using 10 I/O pins, an insufficient length for many cryptographic applications. Second, the PUF evaluation metrics, particularly the inter-device Hamming distance, fell below ideal values. Moreover, the raw responses exhibited statistical bias, rendering them unsuitable for direct use as cryptographic keys. Third, the performance evaluation was limited to Hamming distance analysis.
@@
This section reviews works related to ioPUF+, focusing on PUFs that exploit I/O characteristics and highlighting how they differ from our approach. Furthermore, we examine literature on PUF-based cryptographic key generation methods and compare them with ioPUF+ implementation.
II-A PUFs Related to IOs
@@
The study by Basak et al. measures input currents corresponding to externally applied DC voltages at various IO pins of an IC [15]. Since the intrinsic resistances of pins vary across devices, the pins draw varying currents from the voltage source. The authors use these current measurements to create a unique device identifier for the IC. Our approach differs from that of Basak et al. in the following ways. First, unlike Basak et al., who focus on the measurement of intrinsic resistances of pins, we measure the resistances of the pull-up and pull-down resistors associated with the IOs. Since the pull-up/down resistor values are significant compared to intrinsic pin resistances, the measurement apparatus required is less involved. Second, In ioPUF+, the resistance measurements and generation of the PUF IDs are executed within the IC, eliminating the need for any external measurement apparatus. In contrast, the method in Basak et al. requires external voltage sources and high-precision ammeters connected to different pins of the SoC for input pin current or pin resistance measurements. Their method is less suited for PUF key generation in IoT nodes, which typically cannot accommodate costly, bulky measuring instruments. Therefore, Basak et al.’s method is more suited for lab-based device identification than real-world IoT node identifications.
@@
Vaidya et al. explores a PUF-based device identification scheme using IO pins [9]. In their approach, IO pins are driven to logic-low, and the difference in voltages between pairs of pins is measured. For the measurements, they connected IO pins to a differential ADC in the IC through PCB traces external to the IC. Ideally, the ADC reading should be zero, considering both IO pins are at a logic-low state. However, due to the parasitics of the IO pin driver, the PCB traces, and the non-idealities of the ADC, the measurements will yield non-zero results that vary across the devices. Vaidya et al. utilized 32 pairs of IO pins to formulate the PUF ID. Our work significantly differs from Vaidya et al. in several ways. First, our PUF relies on the pull-up/down resistors on the IO pins, which are in place by design. In contrast, the PUF by Vaidya et al. relies on parasitics and non-idealities, elements not intentionally designed into the system. Reliance on parasitics for generating PUF keys presents several challenges. First, the parasitics are highly dependent on the device’s operation. For example, the ground bounce effect significantly influences the logic low voltage on a device’s IO pins. When the device consumes significant current during its operation, the ground bounce effect kicks in, which can cause the IO pin voltage to increase, thereby affecting the PUF keys. While the effects like ground bounce may also impact the measurement of pull-up/down resistors, the impact is substantially less due to their higher resistance values (several kiloohms) compared to the parasitic resistances (a few milliohms). Further, the characteristics of pull-up/down resistor values and their variations are thoroughly understood and documented by IC manufacturers, facilitating more reliable PUF design. Additionally, unlike our method, Vaidya et al. requires IO pins to be dedicated solely to PUF generation.
@@
Vinko et al. [16] present an analog RC PUF. The method uses the internal pull-ups of two IO pins to charge two external capacitors, which are subsequently discharged through external resistors. The PUF response is obtained from measuring the discharge time of these two capacitors. In comparison to our work, the design by Vinko et al. has several shortcomings. Due to the limited number of IO pins used, the analog RC PUF generates relatively short PUF responses, limiting its applicability in cryptographic applications such as secret-key generation. Vinko et al. average multiple discharge-time measurements to obtain a reliable PUF response, which significantly increases the PUF response generation time. Furthermore, in their design, each IO pin used for PUF measurements requires a dedicated external RC connection; therefore, the pin cannot be used for any other microcontroller function. In contrast, in our work, no passive components are connected to the pins whose pull-up/down resistors are measured. Finally, unlike our approach, the study by Vinko et al. does not present detailed performance metrics or describe the process of generating PUF IDs from raw PUF responses.
@@ II-B Secret Keys Using PUFs
Maes et al. [19] present PUFKY, which is a PUF-based key generator that employs an optimized ring oscillator PUF with Lehmer-Gray encoding. PUFKY achieves high entropy density through a three-step process involving frequency normalization to remove bias, Lehmer encoding for stable ordering representation, and XOR-based entropy compression. This approach, however, increases implementation complexity, utilizing up to 1024 oscillators, concatenated error-correction codes, and a custom coprocessor for BCH decoding. PUFKY generates 128-bit keys with a resource usage of 17% on FPGA. In contrast, our approach utilizes only 6.4% of resources on a low-cost, resource-constrained microcontroller.
@@ III-A I/O-Based PUF Primitive
Pull-up and pull-down resistors are present on the I/O pins of various ICs. Many of these ICs allow attaching/detaching of these resistors through firmware-based register writes. These resistors enable I/O pins for communication and logic operations. For instance, I2C-based communication requires pull-up resistor support for I/O pins [20]. Figure 1 a) shows the I/O architecture of a microcontroller IC, PSoC-5, from Infineon Technologies [21]. In the figure, and represent the pull-up and pull-down resistors, respectively. These resistors can be attached/detached from the I/O by turning transistors and ON or OFF. PSoC-5 allows these transistors to be turned ON/OFF through firmware. The resistance values of these internal resistors vary between pins and across devices due to manufacturing process variations. ioPUF+ utilizes these variations to produce unique PUF responses for the ICs.
@@ III-C Derivation of PUF ID and Secret Key
The regeneration phase is executed whenever the PUF ID is required. In this phase, a fresh PUF response, R”, is recorded. This response may differ from the from the enrolled ID by a small number of bits due to noise or environmental variations. Using the stored helper data, the BCH decoder corrects these errors and reconstructs the original PUF ID.
@@
Although the helper data is stored in non-volatile memory and may be accessible to an attacker, it alone does not reveal information about the PUF responses or the PUF ID. The helper data contains only the information necessary for the BCH decoder to correct errors in PUF responses. Even if an adversary obtains the helper data, they cannot reconstruct the original PUF ID without access to the PUF responses, which are generated only at runtime and are never stored in non-volatile memory. Furthermore, the helper data is a many-to-one function of the PUF response, meaning that multiple distinct responses can yield the same helper data. Therefore, reverse-engineering PUF responses or the PUF ID from the helper data of a device is computationally difficult.

### v2

@@ Abstract
In this work, we present ioPUF+, which incorporates a novel Physical Unclonable Function (PUF) that generates unique fingerprints for Integrated Circuits (ICs) and the IoT nodes encompassing them. The proposed PUF generates device-specific responses by measuring the pull-up and pull-down resistor values on the I/O pins of the ICs, which naturally vary across chips due to manufacturing-induced process variations. Since these resistors are already integrated into the I/O structures of most ICs, ioPUF+ requires no custom circuitry and no new IC fabrication. This makes ioPUF+ suitable for cost-sensitive embedded systems built from Commercial Off-The-Shelf (COTS) components. Beyond introducing a new PUF, ioPUF+ includes a complete datapath for converting raw PUF responses into cryptographically usable secret keys using BCH error correction and SHA-256 hashing. Furthermore, ioPUF+ demonstrates a practical use case of PUF-derived secret keys in securing device-to-device communication using AES encryption.
We implemented ioPUF+ on the Infineon PSoC-5 microcontroller and evaluated its performance across 32 devices using standard PUF metrics. The results show excellent reliability of 100% (zero intra-device Hamming distance), strong uniqueness (inter-device Hamming distance of 50.33%), near-ideal uniformity (50.54%), and negligible bit aliasing. Stability tests under temperature and supply-voltage variations show worst-case bit-error rates of only 2.63% and 2.10%, respectively. We also profiled the resource and energy usage of the complete ioPUF+ system, including the PUF primitive, BCH decoding, SHA-256 hashing, and AES encryption. The full implementation requires only 19.8 KB of Flash, exhibits a latency of 600 ms, and consumes 79 mW of power, demonstrating the suitability of ioPUF+ for resource-constrained IoT nodes.
@@ I Introduction
In many Internet of Things (IoT) applications, securing data is paramount. Consider the case of wearable sensors that continuously collect and transmit sensitive health information from a patient to a nearby hub over the air. If the transmitted data is not adequately protected, an adversary in the vicinity may eavesdrop on the wireless channel as illustrated in Figure 1(a), gaining access to private health information, and potentially misuse it [2].
@@
A common approach to secure data in transit or at rest is to encrypt it using a secret key, often stored in non-volatile memory of the device, following well-established standards such as AES. However, this method has a critical limitation: if an attacker gains physical access to the device as shown in Figure 1(b), the stored key can be extracted [3]. Once compromised, the adversary can eavesdrop on future encrypted communications or decrypt any stored information.
Public-key cryptography offers a solution, in particular for protecting data during transit. Here, the recipient’s public key is provisioned to the IoT node during an enrollment phase. Later, the node encrypts its data using the public key before transmitting it over the air. Only the intended recipient, who owns the corresponding private key, can decrypt it. However, this method has two major drawbacks. First, public-key cryptography incurs significantly higher computational and energy overheads compared to symmetric-key methods like AES, making it less suitable for resource and energy-constrained IoT nodes [4]. Second, if an attacker learns the data format, they may create a malicious node that encrypts forged data using the same public key. The receiver may then process this malicious data as legitimate, leading to data-integrity violations [3]. In safety-critical systems, such as in industrial IoT systems, such violations may trigger catastrophic failures or prolonged downtime.
Another possibility is to protect unauthorized readouts from the device’s memory by storing the key. The method works by disabling the readout circuit to the memory once the key is logged in, so that external devices, e.g., programmers, cannot perform any memory reads from the device. However, the method cannot prevent sophisticated invasive attacks, such as the case where the chip is decapsulated, and the internal circuitry is probed to infer the key.
@@
As an alternative to the aforementioned traditional approaches, Physically Unclonable Function (PUF)-based methods for securing IoT nodes have been proposed. PUF-based techniques eliminate the need to store keys in non-volatile memory [6, 7, 8], thereby reducing the risk of key extraction and the overhead of protecting stored keys. Instead, they generate secret keys during the runtime by exploiting intrinsic hardware characteristics unique to each device, making these keys extremely difficult to duplicate as depicted in Figure 1(c). The generated key resides only in volatile memory (e.g., RAM). As such, when the device powers off, the key disappears from the memory, preventing an attacker from extracting it through physical access. Furthermore, because the key exists only during the powered-on state, additional layers of protection, such as active tamper-detection sensors or probe detection circuits, which operate only when the device is powered, can be employed.
@@
At the core of any PUF-based method for generating secret keys lies a PUF primitive, a circuit and/or algorithm that produces unique, device-specific responses that serve as a fingerprint of the hardware. Several types of PUF primitives have been reported in the literature. However, many of them suffer from a significant drawback: they require custom circuitry and, consequently, a dedicated Integrated Circuit (IC) [9]. For example, the Ring Oscillator PUF (RO-PUF) [10] relies on a specially engineered ring-oscillator network. As such, many of the existing PUF primitives [11, 12, 13] have limited application in IoT nodes, which are typically built using low-cost, resource-constrained Commercial Off-The-Shelf (COTS) components as illustrated in Figure 1(d). Memory-based PUFs, such as SRAM PUFs [14], are compatible with COTS microcontrollers but introduce their own constraints: response generation is typically restricted to the device’s boot phase. As a result, the derived secret key must be stored in memory throughout the device’s runtime, which may raise security concerns.
@@
In response to these limitations, we introduce ioPUF+, which incorporates a novel PUF primitive that generates unique fingerprints for Integrated Circuits (ICs), specifically, microcontroller ICs commonly used in IoT nodes, by measuring the values of the pull-up and pull-down resistors connected to the IC’s input-output (I/O) pins as demonstrated in Figure 1(e). Our approach is based on the observation that, in many IC designs, the pull-up/down resistor values exhibit substantial inter-device variation while showing much smaller intra-device variation across repeated measurements. By leveraging these resistor values, the proposed PUF primitive produces device-specific responses without requiring custom circuitry or new IC fabrication, as pull-up and pull-down resistors are already integral to the IO structures of most COTS microcontroller ICs. Moreover, unlike SRAM PUFs, which restrict response generation to the boot phase, the responses here can be generated at any point during device operation. Consequently, ioPUF+ based secret keys can be derived on a need basis and can be securely erased from memory after use, improving the security of the keys.
@@
The rest of the paper is organized as follows: Section II reviews related work on I/O-based PUFs and PUFs for secret key generation used in cryptography applications. Section III presents the system design of our ioPUF+, the algorithm used for generating ioPUF+ responses, and the procedure for obtaining the PUF ID and secret key from the raw ioPUF+ response. Section IV describes the experimental setup, including hardware implementation on the PSoC-5 microcontroller, evaluation of standard PUF metrics, and discussion of ioPUF+ results for these metrics. In Section V, we demonstrate cryptographic key generation, provide an end-to-end secure communication application, and discuss resource utilization for all components on the PSoC-5 microcontroller. Section VI concludes the paper with future directions.
@@ II Related Work
However, the ioPUF [1] design had several shortcomings. First, it measured either the pull-up or the pull-down resistor of each I/O pin, limiting the PUF response length to 45 bits when using 10 I/O pins, an insufficient length for many cryptographic applications. Second, the PUF evaluation metrics, particularly the inter-device Hamming distance, fell below ideal values. Moreover, the raw responses exhibited statistical bias, rendering them unsuitable for direct use as cryptographic keys. Third, the performance evaluation was limited to Hamming distance analysis.
@@
This section reviews works related to ioPUF+, focusing on PUFs that exploit I/O characteristics and highlighting how they differ from our approach. Furthermore, we examine literature on PUF-based cryptographic key generation methods and compare them with the ioPUF+ implementation.
II-A PUFs Related to I/Os
@@
The study by Basak et al. measures input currents corresponding to externally applied DC voltages at various I/O pins of an IC [15]. Since the intrinsic resistances of pins vary across devices, the pins draw varying currents from the voltage source. The authors use these current measurements to create a unique device identifier for the IC. Our approach differs from that of Basak et al. in the following ways. First, unlike Basak et al., who focus on the measurement of intrinsic resistances of pins, we measure the resistances of the pull-up and pull-down resistors associated with the I/Os. Since the pull-up/down resistor values are significant compared to intrinsic pin resistances, the measurement apparatus required is less involved. Second, in ioPUF+, the resistance measurements and generation of the PUF IDs are executed within the IC, eliminating the need for any external measurement apparatus. In contrast, the method in Basak et al. requires external voltage sources and high-precision ammeters connected to different pins of the SoC for input pin current or pin resistance measurements. Their method is less suited for PUF key generation in IoT nodes, which typically cannot accommodate costly, bulky measuring instruments. Therefore, Basak et al.’s method is more suited for lab-based device identification than real-world IoT node identification.
@@
Vaidya et al. explores a PUF-based device identification scheme using I/O pins [9]. In their approach, I/O pins are driven to logic-low, and the difference in voltages between pairs of pins is measured. For the measurements, they connected I/O pins to a differential ADC in the IC through PCB traces external to the IC. Ideally, the ADC reading should be zero, considering both I/O pins are at a logic-low state. However, due to the parasitics of the I/O pin driver, the PCB traces, and the non-idealities of the ADC, the measurements will yield non-zero results that vary across the devices. Vaidya et al. utilized 32 pairs of I/O pins to formulate the PUF ID. Our work significantly differs from Vaidya et al. in several ways. First, our PUF relies on the pull-up/down resistors on the I/O pins, which are in place by design. In contrast, the PUF by Vaidya et al. relies on parasitics and non-idealities, elements not intentionally designed into the system. Reliance on parasitics for generating PUF keys presents several challenges. First, the parasitics are highly dependent on the device’s operation. For example, the ground bounce effect significantly influences the logic low voltage on a device’s I/O pins. When the device consumes significant current during its operation, the ground bounce effect kicks in, which can cause the I/O pin voltage to increase, thereby affecting the PUF keys. While the effects like ground bounce may also impact the measurement of pull-up/down resistors, the impact is substantially less due to their higher resistance values (several kiloohms) compared to the parasitic resistances (a few milliohms). Further, the characteristics of pull-up/down resistor values and their variations are thoroughly understood and documented by IC manufacturers, facilitating more reliable PUF design. Additionally, unlike our method, Vaidya et al. requires I/O pins to be dedicated solely to PUF generation.
@@
Vinko et al. [16] present an analog RC PUF. The method uses the internal pull-ups of two I/O pins to charge two external capacitors, which are subsequently discharged through external resistors. The PUF response is obtained from measuring the discharge time of these two capacitors. In comparison to our work, the design by Vinko et al. has several shortcomings. Due to the limited number of I/O pins used, the analog RC PUF generates relatively short PUF responses, limiting its applicability in cryptographic applications such as secret-key generation. Vinko et al. average multiple discharge-time measurements to obtain a reliable PUF response, which significantly increases the PUF response generation time. Furthermore, in their design, each I/O pin used for PUF measurements requires a dedicated external RC connection; therefore, the pin cannot be used for any other microcontroller function. In contrast, in our work, no passive components are connected to the pins whose pull-up/down resistors are measured. Finally, unlike our approach, the study by Vinko et al. does not present detailed performance metrics or describe the process of generating PUF IDs from raw PUF responses.
@@ II-B Secret Keys Using PUFs
Maes et al. [19] present PUFKY, which is a PUF-based key generator that employs an optimized ring oscillator PUF with Lehmer-Gray encoding. PUFKY achieves high entropy density through a three-step process involving frequency normalization to remove bias, Lehmer encoding for stable ordering representation, and XOR-based entropy compression. This approach, however, increases implementation complexity, utilizing up to 1024 oscillators, concatenated error-correction codes, and a custom coprocessor for BCH decoding. PUFKY generates 128-bit keys with a resource usage of 17% on an FPGA. In contrast, our approach utilizes only 6.4% of resources on a low-cost, resource-constrained microcontroller.
@@ III-A I/O-Based PUF Primitive
Pull-up and pull-down resistors are present on the I/O pins of various ICs. Many of these ICs allow attaching/detaching of these resistors through firmware-based register writes. These resistors enable I/O pins for communication and logic operations. For instance, I2C-based communication requires pull-up resistor support for I/O pins [20]. Figure 2 (a) shows the I/O architecture of a microcontroller IC, PSoC-5, from Infineon Technologies [21]. In the figure, and represent the pull-up and pull-down resistors, respectively. These resistors can be attached/detached from the I/O by turning transistors and ON or OFF. PSoC-5 allows these transistors to be turned ON/OFF through firmware. The resistance values of these internal resistors vary between pins and across devices due to manufacturing process variations. ioPUF+ utilizes these variations to produce unique PUF responses for the ICs.
@@ III-C Derivation of PUF ID and Secret Key
The regeneration phase is executed whenever the PUF ID is required. In this phase, a fresh PUF response, R”, is recorded. This response may differ from the enrolled ID by a small number of bits due to noise or environmental variations. Using the stored helper data, the BCH decoder corrects these errors and reconstructs the original PUF ID.
@@
Although the helper data is stored in non-volatile memory and may be accessible to an attacker, it alone does not reveal information about the PUF responses or the PUF ID. The helper data contains only the information necessary for the BCH decoder to correct errors in PUF responses. Even if an adversary obtains the helper data, they cannot reconstruct the original PUF ID without access to the PUF responses, which are generated only at runtime and are never stored in non-volatile memory. Furthermore, the helper data is a many-to-one function of the PUF response, meaning that multiple distinct responses can yield the same helper data [24]. Therefore, reverse-engineering PUF responses or the PUF ID from the helper data of a device is computationally difficult.